find the consecutive odd itergers such that twice the larger is 17 more than the smaller
Answer by checkley77(7072)

LET X & X+2 BE THE 2 ODD INTEGERS.
2(X+2)=X+17
2X+4=X+17
2X-X=17-4
X=13 ANS.
13+2=15 ANS.
PROOF:
2*15=13+17
30=30